Marginal tax rates   Partner​ A, a single​ taxpayer, is one of two partners in a small business. As​ such, she receives​ pass-through income that is taxed at her personal tax rates. After all adjustments and deductions have been​ made, including the​ 20% qualified business income​ deduction, she is preparing to calculate her taxes owed for the year. Using the tax rate schedule given here (SEE BELOW) perform the​ following:

a. Find the marginal tax rate for the following levels of sole proprietorship earnings before​ taxes:

$16500, $58200, $91500, $151000, $251000, $452000 and $1.4 million

b. Plot the marginal tax rates​ (measured on the ​y-​axis) against the pretax income levels​ (measured on the ​x-​axis). Explain the relationship between these variables.

Tax Rate - Single Taxpayers -2018

Tax calculation Range of taxable income Base tax + (Marginal rate × amount over base bracket)

$0 to $9,525 . $0 + (10% × amount over $0)

9,525 to 38,700 953 + (12% × amount over 9,525)

38,700 to 82,500 . 4,454 + (22% × amount over 38,700)

82,500 to 157,500 14,090 + (24% × amount over 82,500)

157,500 to 200,000 32,090 + (32% × amount over 157,500)

200,000 to 500,000 45,690 + (35% × amount over 200,000)

Over 500,000 150,690 + (37% × amount over 500,000)

a. The marginal tax rate for earnings before taxes of $16,500 is .....​%. ​(Round to the nearest​ integer.)
Answer #1

(a) Statement Showing Marginal Tax rates at differnent level of incomes

Income befor taxes tax rate on base income tax on excess income total tax Marginal tax rate
$16500 upto 9525=953

16500-9525=6975*12%

=837

=953+837

=1790

12%
$58200 upto 38700=4454

58200-38700=19500*22%

=4290

=4454+4290

=8744

22%
$91500 upto 82500=14090

91500-82500=9000* 24%

=2160

=14090+2160

=16250

24%

$151000

upto 82500=14090

151000-82500=68500

=68500*24%

=16440

=16440+14090

=30530

24%
$251000 upto 200000 =45690

251000-200000=51000

=51000*35%

=17850

=45690+17850

=63540

35%
$452000 upto 200000=45690

452000-200000=252000

=252000*35%

=88200

=45690+88200

=133890

35%
$1400000 upto 500000=150690

1400000-500000=900000

=900000*37%

=333000

=333000+150690

=483690

37%

Marginal Tax rate is rate of tax which is paid on higher level of income which means every excess income over base income paid with the highest rate of tax .

(B) Statement of graph with there relationship shown in attachment

as lower income $9525 shows marginal tax rate to be 10% and every single excess money will be charged @ 12% marginal tax upto income 38700 and when income above 500000 is charged @ 37%